# Retrosynthesis Planning via Worst-path Policy Optimisation in Tree-structured MDPs

### v1

@@ 3.3 Value Estimation
The advantage function quantifies the relative benefit of applying reaction $a$ to molecule $s$ compared to following policy $\pi$:
@@
A positive advantage indicates that reaction $a$ leads to better outcomes than the policy’s average behaviour. The key insight of our method is to prioritise imitation of past decisions with high advantages. The following proposition establishes a relationship between Q-value and value functions, enabling more efficient advantage estimation:
@@
The Q-value function $Q^{\pi}(s,a)$ equals its immediate reward plus the discounted value of its next states:
$Q^{\pi}(s,a)=r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime})$ (6)
The proof is provided in Appendix C.1.
Based on this proposition, we can estimate the advantage as:
@@
Furthermore, Proposition 2 shows that $V^{\pi}(s)$ satisfies a recursive relation, providing the basis for value function approximation:
@@
The proof is provided in Appendix C.2.
@@
The proof is provided in Appendix C.3.
@@ 3.4 Self-imitation Learning for Retrosynthesis
A key challenge in retrosynthesis is ensuring that the learnt policy proposes realistic chemical reactions at each step. To accomplish this, we leverage a pretrained single-step model $\pi^{0}$ that reflects expert or empirical knowledge from the training set. Single-step models empirically produce valid chemical transformations with high fidelity, motivating us to maintain their support when learning an improved policy. Specifically, we define the constrained policy set:
@@
Our objective is to find a policy $\pi\in\Pi$ that maximises the worst-path objective. We achieve this through an iterative procedure. In each iteration $i$, we aim to find an improved policy $\pi^{i+1}$ by imitating advantageous state-action pairs $(s,a)$ experienced under policy $\pi^{i}$. The advantage $A^{\pi^{i}}(s,a)$ quantifies this, and the learning objective for $\pi^{i+1}$ is formulated as:
@@
where $\beta>0$ is the advantage coefficient controlling the strength of advantage weighting, and $d_{\pi^{i}}$ is the state distribution induced by policy $\pi^{i}$ [29]. Intuitively, reactions with higher advantages receive higher weights, guiding the policy toward better-than-average reactions.
Because each new policy $\pi^{i+1}$ is derived by re-weighting $\pi^{i}$, and the initial policy $\pi^{0}$ restricts the support, the entire policy sequence $\{\pi^{i}\}_{i\geq 0}$ remains within the feasible set $\Pi$ [12]. This ensures that all proposed reactions throughout training remain chemically valid.
We now show that this self-imitation update leads to monotonic improvement in the worst-path value function:
@@
The proof is provided in Appendix C.4.
@@
This update re-weights the previous policy by exponentiated advantages, increasing the likelihood of actions that outperform the current expectation.
For policy optimisation, we use the current policy $\pi^{i}$ to interact with the retrosynthesis environment and collect data to learn the next iteration’s policy $\pi^{i+1}$. Through iterative weighted imitation, the policy increasingly favours high-quality reactions that lead to successful synthetic routes with shorter paths, while maintaining chemical plausibility by respecting the support of the pretrained model.
@@ 4 The InterRetro Algorithm
0: Pretrained one-step policy $\pi_{\theta}$, value function $V_{\phi}$, training dataset $\mathcal{D}$, replay buffer $\mathcal{B}$.
@@
We now present our algorithm InterRetro for retrosynthesis planning. Section 4.1 introduces how the single-step model interacts with the tree MDP and collects experiences. Section 4.2 describes how to learn a value function on the worst-path objective and how to fine-tune the policy to reproduce advantageous decisions. Finally, Section 4.3 provides more implementation details for reproducibility.
@@ 4.1 Environment Interactions
The Explore procedure in Algorithm 1 shows how InterRetro constructs a synthesis tree by interacting with the tree MDP. Starting from a target molecule $m$, the single-step model proposes a reaction $a$ to decompose the molecule into a set of reactants $\mathcal{S}_{r}$. These reactants will be attached to the parent node $m$ in the synthesis tree. Among them, the non-building blocks are then placed in a collection (e.g., a queue) for further expansion. Each subsequent round pops a molecule from the collection to continue the decomposition process. The interaction terminates when the collection becomes empty (meaning all leaf nodes are building blocks) or when a predefined maximum number of steps is reached.
@@ 4.2 Value Function and Policy Learning
The Explore procedure constructs synthesis trees by interacting with the tree MDP. We extract all branches $(s,a,\mathcal{S}_{r})$ within successful subtrees and store them in a first-in first-out (FIFO) replay buffer $\mathcal{B}$ for value function and policy learning (see InterRetro in Algorithm 1).
@@
$\mathcal{L}(\phi)=\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}\left[(V_{\phi}(s)-(r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{S}_{r}}V_{\phi^{-}}(s^{\prime})))^{2}\right],$ (14)
@@
The policy network $\pi_{\theta}$ with parameter $\theta$ is updated using the weighted imitation learning objective from Eq. 11, implemented as the following loss function:
$\mathcal{L}(\theta)=\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}\left[\exp_{\mathrm{clip}}(\beta A_{\phi}(s,a))\log\pi_{\theta}(a\mid s)\right],$ (15)
where $\beta>0$ is the advantage coefficient which controls the imitation strength on the past successful experience, and $\exp_{\mathrm{clip}}(\cdot)$ is the exponential function with a clipped output range $(0,C]$ for numerical stability. The advantage is estimated by the value function network, according to Proposition 1:
@@
This joint optimisation of the value and policy networks enables InterRetro to iteratively improve its planning policy while maintaining robustness and chemical validity.
@@ 4.3 Implementation Details
We choose Graph2Edits as the single-step model due to its strong performance and flexibility in single-step retrosynthesis prediction [38]. Graph2Edits represents a molecule through a graph and predicts a sequence of graph edits to transform it to reactants. These graph edits can delete bonds, modify bond types, alter atoms, or attach leaving groups (see Appendix D for more details). The single-step model and our proposed value function encode the molecule using a message passing neural network [33] and forward the latent code into linear layers with ReLU as the activation function. For training, we run $6$ parallel exploration processes and collect $36$ synthetic trees per iteration. The networks are updated $5$ times per round using data from a compact replay buffer of $20,000$ branches, chosen to reduce CPU memory usage and maintain close alignment between the data-collecting policy $\pi^{i}$ and the updated policy $\pi^{i+1}$. Our models are trained on a single NVIDIA RTX A5000 GPU and, without pre-training the single-step model, require approximately $48$ hours to fully converge. The code will be open-sourced after the publication of this manuscript.
@@ 5 Experimental Results
In this section, we aim to answer the following questions: (1) What are the advantages of our proposed method compared to the SOTA algorithms? (2) How does each component of the method contribute to the performance? Additionally, we examine the real-world feasibility of the proposed synthetic routes and present illustrative examples in Appendix B.
@@ 5.1 Benchmark Results
The proposed method is trained by creating synthetic routes for nearly $300k$ molecules in the USPTO-50k dataset with commercially available building blocks from the eMolecules dataset111eMolecules: https://downloads.emolecules.com/free/.. We evaluate performance on three benchmarks of increasing difficulty: Retro*-190 [1], ChEMBL-1000, and GDB17-1000 [10], where the suffix indicates the dataset size. ChEMBL-1000 from ChEMBL dataset [35] is a manually curated collection of bioactive molecules exhibiting drug-like properties, while GDB17-1000 is a subset of the GDB-17 dataset [18], which covers a size range common for typical drugs and lead compounds. We compare against established retrosynthesis planning methods: MCTS, Retro* [1], Self-improve [9], PDVN [10], and GraphRetro [31]. Additionally, we include two single-step methods, MEGAN [19] and Graph2Edits [38], combined with Retro* as search algorithm, and two recently proposed methods, DreamRetroer [36] and RetroCaptioner [11]. Baseline results are produced from their official implementations or the Syntheseus project [13].
Success Rate.
@@
Success rate measures the percentage of target molecules that can be successfully decomposed into building blocks. In Table 1, we firstly compare methods under different model-call budgets: $100$, $200$, and $500$. Our proposed InterRetro significantly outperforms SOTA algorithms across all three test sets. With $500$ model calls, InterRetro achieves $100\%$, $98.2\%$, and $99.5\%$ success rates on Retro*-190, ChEMBL-1000, and GDB17-1000, respectively. All synthetic routes generated by InterRetro on Retro*-190 using 500 model calls are provided in the supplementary materials.
@@
Table 1: Performance evaluation on three benchmarks. The evaluation metrics include the success rate under different test molecules with different budget of model calls, which are direct generation (DG), $100$, $200$ and $500$ model calls. The DG columns are single-step model’s results without search.
@@
Retro-190 ChEMBL-1000 GDB17-1000 Single-step Search DG 100 200 500 DG 100 200 500 DG 100 200 500 Template MCTS $20.00$ $43.68$ $47.37$ $62.63$ $32.00$ $45.60$ $68.80$ $71.90$ $3.00$ $3.20$ $3.70$ $4.50$ Template Retro* $20.00$ $38.42$ $58.42$ $79.47$ $32.00$ $69.10$ $72.00$ $74.70$ $3.00$ $5.40$ $6.60$ $7.50$ LocalRetro MCTS 22.10 $44.21$ $57.36$ $62.10$ $47.30$ $62.70$ $69.10$ $75.00$ $4.60$ $14.00$ $16.70$ $20.30$ LocalRetro Retro* $22.10$ $58.94$ $64.73$ $73.68$ $47.30$ $74.80$ $80.40$ $82.40$ $4.60$ $18.90$ $22.20$ $28.80$ MEGAN Retro* $8.42$ $60.52$ $62.10$ $73.15$ $38.00$ $71.70$ $75.40$ $79.00$ $6.20$ $37.60$ $45.70$ $57.20$ Graph2Edits Retro* $16.84$ $41.05$ $50.00$ $56.31$ $47.10$ $68.70$ $78.80$ $80.70$ $5.90$ $18.20$ $24.00$ $32.20$ Self-improve Retro* - $67.37$ $83.16$ $96.32$ - - - $81.10$ - - - $15.00$ PDVN Retro* - $93.68$ $97.37$ $98.95$ - - - $83.50$ - - - $26.90$ RetroCaptioner Retro* $5.26$ $68.94$ $72.63$ $85.26$ $3.90$ $72.60$ $76.50$ $78.70$ $3.20$ $56.20$ $68.20$ $75.20$ DreamRetroer $32.10$ $78.94$ $88.42$ $90.52$ $31.10$ $78.10$ $81.70$ $83.10$ $4.20$ $27.36$ $28.97$ $33.20$ InterRetro MCTS $\boldsymbol{95.78}$ $89.47$ $98.94$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $78.40$ $89.30$ $97.50$ $\boldsymbol{89.00}$ $80.80$ $96.10$ 99.50 InterRetro Retro* $\boldsymbol{95.78}$ $\boldsymbol{96.31}$ $\boldsymbol{100.00}$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $\boldsymbol{91.40}$ $\boldsymbol{96.20}$ $\boldsymbol{98.20}$ $\boldsymbol{89.00}$ $\boldsymbol{83.80}$ $\boldsymbol{96.50}$ $97.20$
@@
Route Length.
Route length is the number of reactions needed to synthesise a target molecule. Route length could be a concern with our worst-path objective since it does not consider the width of the tree and the total number of reactions required. In Table 2, we compare the route length with other baselines on the $138$ target molecules that all methods can resolve. Our method outperforms the SOTA by $4.85\%$.
Figure 4: Experimental figures. (a) Performance under different training data usage and computation budgets. (b) Statistics on estimated depth of synthesis trees. (c) Ablations on the advantage coefficient.
Sample Efficiency.
Sample efficiency refers to the amount of training data required to learn an effective policy. We evaluate the policy using subsets of the training data at $1\%$, $2\%$, $5\%$, $10\%$, and $100\%$, corresponding to approximately $3k$, $6k$, $15k$, $30k$, and $300k$ molecules. As shown in Figure 4(a), InterRetro achieves near SOTA performance ($92\%$) with just $10\%$ of the training set when directly generating synthesis trees, and reaches SOTA performance when combined with $200$ Retro* search iterations. Notably, when trained on the full dataset, InterRetro outperforms most existing methods in direct generation, achieving a success rate of $95.78\%$. Similar trends are observed on the ChEMBL-1000 and GDB17-1000 benchmarks.

### v2

@@ 3.3 Advantage Estimation
With these definitions in place, we can express the advantage function, which quantifies the relative benefit of applying reaction $a$ to molecule $s$ compared to following the policy:
@@
A positive advantage indicates that reaction $a$ leads to better outcomes than the policy’s average behaviour. Proposition 1 establishes a relationship between $Q^{\pi}(s,a)$ and $V^{\pi}(s)$:
@@
The Q-function $Q^{\pi}(s,a)$ equals its immediate reward plus the discounted value of its next states:
$Q^{\pi}(s,a)=r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}).$ (7)
The proof is provided in Appendix B.1.
Thus, we can estimate the advantage as:
@@
Next, we derive a recursive form of the value function, which forms the foundation for learning a parameterised value function.
@@
The proof is provided in Appendix B.2.
@@
The proof is provided in Appendix B.3.
@@ 3.4 Self-imitation Learning for Retrosynthesis
A key challenge in retrosynthesis is ensuring that the learned policy proposes chemically valid reactions at each step. To address this, we constrain policy learning within the support of a pre-trained single-step model, which can empirically capture high-fidelity chemical transformations. Specifically, we leverage a pre-trained single-step model $\pi^{0}$, trained to reflect expert or empirical reaction knowledge from the dataset zhong2023graph2edits, and define the constrained policy set:
@@
Our objective is to find a policy $\pi\in\Pi$ that maximises the worst-path objective. We achieve this through an iterative procedure: in each iteration $i$, we aim to find an improved policy $\pi^{i+1}$ by imitating advantageous state-action pairs $(s,a)$ experienced under policy $\pi^{i}$. The advantage $A^{\pi^{i}}(s,a)$ quantifies this, and the learning objective for $\pi^{i+1}$ is formulated as:
@@
where $\beta>0$ is the advantage coefficient controlling the strength of advantage weighting, and $d_{\pi^{i}}$ is the state distribution induced by policy $\pi^{i}$ wang2018marwil. In this case, reactions with higher advantages receive higher weights, guiding the policy toward better-than-average reactions.
Since each new policy $\pi^{i+1}$ is derived by re-weighting $\pi^{i}$, and the initial policy $\pi^{0}$ restricts the support, the entire policy sequence $\{\pi^{i}\}_{i\geq 0}$ remains within the feasible set $\Pi$ mao2023str. This ensures that all proposed reactions throughout training remain chemically valid.
@@
The proof is provided in Appendix B.4.
@@
which re-weights the previous policy by the exponentiated advantages, thereby increasing the probability of actions that yield higher returns than the current expectation.
For policy optimisation, we use the current policy $\pi^{i}$ to interact with the retrosynthesis environment and collect data to learn the next iteration’s policy $\pi^{i+1}$. Through iterative weighted imitation, the policy increasingly favours high-quality reactions that lead to successful synthetic routes with shorter paths, while maintaining chemical plausibility by respecting the support of the pre-trained model.
@@ 4 The InterRetro Algorithm
0: pre-trained one-step policy $\pi_{\theta}$, value function $V_{\phi}$, training set $\mathcal{D}$, replay buffer $\mathcal{B}$.
@@
We now present our algorithm InterRetro for retrosynthesis planning. Section 4.1 introduces how the single-step model interacts with the tree MDP and collects trajectories. Section 4.2 describes how to learn a value function on the worst-path objective and how to fine-tune the policy to reproduce advantageous decisions. Finally, Section 4.3 provides more implementation details for reproducibility.
@@ 4.1 Environment Interactions
The Explore procedure in Algorithm 1 shows how InterRetro constructs a synthetic tree by interacting with the tree MDP. Starting from a target molecule $m$, the single-step model proposes a reaction $a$ to decompose the molecule into a set of reactants $\mathcal{S}_{r}$. These reactants will be attached to the parent node $m$ in the synthetic tree. Among them, the non-building blocks are then placed in a collection (e.g., a queue) for further expansion. Each subsequent round pops a molecule from the collection to continue the decomposition process. The interaction terminates when the collection becomes empty (meaning all leaf nodes are building blocks) or when a predefined maximum number of steps is reached.
@@ 4.2 Value Function and Policy Learning
The Explore procedure constructs synthetic trees by interacting with the tree MDP. We extract all branches $(s,a,\mathcal{S}_{r})$ within successful subtrees and store them in a first-in first-out replay buffer $\mathcal{B}$ for value function and policy learning (see InterRetro in Algorithm 1).
@@
$\mathcal{L}(\phi)=\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}[(V_{\phi}(s)-(r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{S}_{r}}V_{\phi^{-}}(s^{\prime})))^{2}],$ (15)
@@
The policy network $\pi_{\theta}$ with pre-trained parameter $\theta$ is updated using the weighted imitation learning objective from Eq. 12, implemented as the following loss function:
$\mathcal{L}(\theta)=-\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}[\exp_{\mathrm{clip}}(\beta A_{\phi}(s,a))\log\pi_{\theta}(a\mid s)],$ (16)
where $\beta>0$ is the advantage coefficient which controls the imitation strength on the past successful experience, and $\exp_{\mathrm{clip}}(\cdot)$ is the exponential function with a clipped output range $(0,C]$ for numerical stability. The advantage is estimated by the value function network, according to Eq. 8:
@@
This joint optimisation of the value and policy networks enables InterRetro to fine-tune the single-step model, increasing the probability of reproducing high-advantage past decisions as indicated by the value function.
@@ 4.3 Implementation Details
We choose Graph2Edits as the single-step model due to its strong performance and flexibility in single-step retrosynthesis prediction zhong2023graph2edits. Graph2Edits represents a molecule through a graph and predicts a sequence of graph edits to transform it to reactants. These graph edits can delete bonds, modify bond types, alter atoms, or attach leaving groups (see Appendix C for more details). The single-step model and our proposed value function encode the molecule using a message passing neural network yang2019dmpnn and forward the latent code into linear layers with ReLU as the activation function. For training, we run $6$ parallel exploration processes and collect $36$ synthetic trees per iteration. The networks are updated $5$ times per iteration using data from a compact replay buffer of maximal $20,000$ branches, chosen to reduce CPU memory usage and maintain close alignment between the data-collecting policy $\pi^{i}$ and the updated policy $\pi^{i+1}$. Our models are trained on a single NVIDIA RTX A5000 GPU and, without pre-training the single-step model, require approximately $48$ hours to fully converge. The code has been open-sourced [^1] GitHub repository: https://github.com/MianchuWang/InterRetro..
@@ 5 Experimental Results
In this section, we aim to answer the following questions: (1) What are the advantages of our proposed method compared to the SOTA algorithms? (2) How does each component of the method contribute to the performance? Additionally, we examine the real-world feasibility of the proposed synthetic routes and present illustrative examples in Appendix A.
@@ 5.1 Benchmark Results
The proposed method is trained by creating synthetic routes for nearly $300k$ molecules in the USPTO-50k dataset with commercially available building blocks from the eMolecules dataset222eMolecules: https://downloads.emolecules.com/free/.. We evaluate performance on three benchmarks of increasing difficulty: Retro*-190 chen2020retrostar, ChEMBL-1000 liu2023pdvn; zdrazil2023chembl, and GDB17-1000 liu2023pdvn; ruddigkeit2012gdb17, where the suffix indicates the dataset size. We compare against established retrosynthesis planning methods: MCTS, Retro* chen2020retrostar, Self-improve kim2021retrostarplus, PDVN liu2023pdvn, and GraphRetro xie2022retrograph. Additionally, we include two single-step methods, MEGAN sacha2021megan and Graph2Edits zhong2023graph2edits, combined with Retro* as the search algorithm, and two recently proposed methods, DreamRetroer zhang2025dreamretroer and RetroCaptioner liu2024retrocaptioner. Baseline results are produced from their official implementations or the Syntheseus project maziarz2025syntheseus.
@@
Success Rate. Success rate measures the percentage of target molecules that can be successfully decomposed into building blocks. In Table 1, we firstly compare methods under different model-call budgets: $100$, $200$, and $500$. Our proposed InterRetro significantly outperforms SOTA algorithms across all three test sets. With $500$ model calls, InterRetro achieves $100\%$, $98.2\%$, and $99.5\%$ success rates on Retro*-190, ChEMBL-1000, and GDB17-1000, respectively. All synthetic routes generated by InterRetro on Retro*-190 using 500 model calls are provided in the supplementary materials.
@@
Table 1: Performance evaluation on three benchmarks. The evaluation metrics include the success rate under different test molecules with different budgets of model calls, which are direct generation (DG), $100$, $200$ and $500$ model calls. The DG columns are single-step model’s results without search.
@@
Retro-190 ChEMBL-1000 GDB17-1000 Single-step Search DG 100 200 500 DG 100 200 500 DG 100 200 500 Template MCTS $20.00$ $43.68$ $47.37$ $62.63$ $32.00$ $45.60$ $68.80$ $71.90$ $3.00$ $3.20$ $3.70$ $4.50$ Template Retro* $20.00$ $38.42$ $58.42$ $75.26$ $32.00$ $69.10$ $72.00$ $74.70$ $3.00$ $5.40$ $6.60$ $7.50$ LocalRetro MCTS 22.10 $44.21$ $57.36$ $62.10$ $47.30$ $62.70$ $69.10$ $75.00$ $4.60$ $14.00$ $16.70$ $20.30$ LocalRetro Retro* $22.10$ $58.94$ $64.73$ $73.68$ $47.30$ $74.80$ $80.40$ $82.40$ $4.60$ $18.90$ $22.20$ $28.80$ MEGAN Retro* $8.42$ $60.52$ $62.10$ $73.15$ $38.00$ $71.70$ $75.40$ $79.00$ $6.20$ $37.60$ $45.70$ $57.20$ Graph2Edits Retro* $16.84$ $41.05$ $50.00$ $56.31$ $47.10$ $68.70$ $78.80$ $80.70$ $5.90$ $18.20$ $24.00$ $32.20$ Self-improve Retro* - $67.37$ $83.16$ $94.74$ - - - $81.10$ - - - $15.00$ PDVN Retro* - $93.68$ $97.37$ $98.95$ - - - $83.50$ - - - $26.90$ RetroCaptioner Retro* $5.26$ $68.94$ $72.63$ $85.26$ $3.90$ $72.60$ $76.50$ $78.70$ $3.20$ $56.20$ $68.20$ $75.20$ DreamRetroer $32.10$ $78.94$ $88.42$ $90.52$ $31.10$ $78.10$ $81.70$ $83.10$ $4.20$ $27.36$ $28.97$ $33.20$ InterRetro MCTS $\boldsymbol{95.78}$ $89.47$ $98.94$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $78.40$ $89.30$ $97.50$ $\boldsymbol{89.00}$ $80.80$ $96.10$ 99.50 InterRetro Retro* $\boldsymbol{95.78}$ $\boldsymbol{96.31}$ $\boldsymbol{100.00}$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $\boldsymbol{91.40}$ $\boldsymbol{96.20}$ $\boldsymbol{98.20}$ $\boldsymbol{89.00}$ $\boldsymbol{83.80}$ $\boldsymbol{96.50}$ $97.20$
@@
Route Length. Route length is the number of reactions needed to synthesise a target molecule. Route length could be a concern with our worst-path objective since it does not consider the width of the tree and the total number of reactions required. In Table 2, we compare the route length with other baselines on the $138$ target molecules that all methods can resolve. Our method outperforms the SOTA by $4.85\%$.
@@
Figure 4: Experimental figures. (a) Performance under different training data usage and computation budgets. (b) Statistics on estimated depth of synthetic trees. (c) Ablations on the advantage coefficient.
Sample Efficiency. Sample efficiency refers to the amount of training data required to learn an effective policy. We evaluate the policy using subsets of the training data at $1\%$, $2\%$, $5\%$, $10\%$, and $100\%$, corresponding to approximately $3k$, $6k$, $15k$, $30k$, and $300k$ molecules. As shown in Figure 4(a), InterRetro achieves near SOTA performance ($92\%$) with just $10\%$ of the training set when directly generating synthetic trees, and reaches SOTA performance when combined with $200$ Retro* search iterations. Notably, when trained on the full dataset, InterRetro outperforms most existing methods in direct generation, achieving a success rate of $95.78\%$. Similar trends are observed on the ChEMBL-1000 and GDB17-1000 benchmarks.